OpenStudy (anonymous):

sally buys 5 shirts on sale for $20 each. She also buys 5 pairs of shorts that all have the same price. The total cost for clothes is $185. How much did each pair of shorts cost?

OpenStudy (mathteacher1729):

What part of this problem is getting you stuck?

OpenStudy (anonymous):

the whole thing

OpenStudy (mathteacher1729):

Ok. So let's start at the beginning. How much does one shirt cost?

OpenStudy (anonymous):

20

OpenStudy (mathteacher1729):

How much do two shirts cost?

OpenStudy (anonymous):

40

OpenStudy (mathteacher1729):

Yup. How much does ... 3 shirts = 4 shirts = 5 shirts = (believe me, I'm going somewhere with this!)

OpenStudy (anonymous):

60 80 100?

OpenStudy (mathteacher1729):

Yes, yes, and yes. :)

OpenStudy (mathteacher1729):

OK so this is what you figured out so far... the 5 shirts cost 100 dollars.

OpenStudy (mathteacher1729):

Now, if the TOTAL money spent is 185 dollars.... how much of that was spent on SHORTS only? (remember, the SHIRTS cost 100 dollars)

OpenStudy (anonymous):

85

OpenStudy (mathteacher1729):

Yup! So 85 dollars on JUST the shorts. But she bought FIVE pairs of shorts... So 85 divided by 5 = ....

OpenStudy (anonymous):

17?

OpenStudy (mathteacher1729):

Yay! :D

OpenStudy (jiteshmeghwal9):

total cost of shirts = 5*20=100$ total cost of shorts=5*x=5x$ A/q, 100$+5x$=185$ 5x$=185$-100$ 5x=85$ x=85/5=17

OpenStudy (anonymous):

so the shorts cost $17 each?

OpenStudy (mathteacher1729):

Yes, you got it @ashleymariel and you did it on your own. :)

OpenStudy (anonymous):

thank you very much (: @mathteacher1729

OpenStudy (mathteacher1729):

You're welcome. :) You can use the same thing we did here for similar problems. Break it down into little bits instead of trying to do everything at once.